Two books each with a mass of 0.9 kg are stacked on a desk. The coefficient of static friction between the books is 0.31 and the coefficient of static friction between the bottom book and the table is 0.24. You yank on the bottom book with a purely horizontal force of 35 N. What is the resulting acceleration of the bottom book?
Answer #1

Normal force N on both books is

N =(m1 + m2) * g = (0.9 + 0.9)*9.8 = 1.8 *9.8

Now as per information given, we have to write equation of motion for the system

F - \mus N = m2 a

therefore, a = (F - \mus N)/ m2

= (35 - (1.8 *9.8))/0.9

= (35-4.2336)/ 0.9

= 30.7664/0.9

= 34.18 m/s2